# Baikal–GVD Collaboration Constraints on the diffuse flux of multi-PeV astrophysical neutrinos obtained with the Baikal Gigaton Volume Detector

### v1

@@ I Introduction
Recently, the KM3Net experiment has reported about the detection of a neutrino with best-fit energy well above 100 PeV [14]. This event could potentially be the first cosmogenic neutrino ever detected [15]. However, to make conclusions from this observation, one should take into account that events of similar energies have not been detected by other neutrino telescopes, despite larger exposure. These non-detections constrain strongly the isotropic neutrino flux, see e.g. [16, 17].
@@ II Data
For the present analysis, we use the Baikal-GVD sample of cascade events detected between Spring 2018 and Spring 2024. In this period, the detector was operated in several, increasingly larger configurations, with one or two new “clusters” of optical modules (OMs) added each spring. One cluster comprises 288 OMs arranged on eight vertical strings, with 36 OMs per string. Over the stated period, the number of clusters increased from three to eleven. The total exposure of the data set used here corresponds to years of one-cluster operation. The reader may find descriptions of the experiment, event reconstruction, analysis techniques and procedures e.g. in Refs. [18, 19, 20, 21, 22].
We select high-energy cascade events with the multiplicity of triggered optical modules (OMs) 8 and impose the quality cuts described in Ref. [19], except for the cut introduced to suppress atmospheric muon bundles, which are not expected to contaminate the sample at these high energies. Relaxing this latter cut and multiplicity of OMs have allowed to increase the exposure considerably.
@@ III Flux limits
where , is the effective area, the integration extends over the bins with in energy and over the data collection time in . The resulting upper limits are listed in Table 1
@@ IV.2 Constraints on the cosmogenic neutrino flux models
Predictions of the selected models are shown in gray color in both Fig. 2 and Fig. 4. To obtain quantitative constraints on the scenarios, we calculate, for each of them, the expected number of events in the KM3-230213A 90% CL energy uncertainty band. Table 2
| best fit for | upper limit | |
| model | cosmogenic | (90% CL) |
| KM3-230213A | ||
| Auger low 2023 [29] | 610 | |
| Auger high 2023 [29] | 11.5 | |
| Muzio maximal 2023 [30] | 0.56 | |
| Ehlert dip 2023 [13] | 0.96 | |
| KM3NeT 2025 extreme [15] | 3.9 |
presents the estimates of model normalization factors obtained assuming that the flux corresponding to a single detected event, KM3-230213A, is cosmogenic. Within this assumption, the lowest-flux Auger 2023 model is disfavored at . We also present there upper limits on the normalization factors, which are free from this assumption.

### v2

@@ I Introduction
However, this cosmogenic neutrino flux is yet to be measured. Only three neutrino events with best-fit reconstructed energies above 5 PeV have been detected by the presently largest neutrino telescope, IceCube. The highest energy of these three is about 13 PeV, see e.g. Ref. [14] for details and further references. Recently, the KM3Net experiment has reported about the detection of a neutrino with best-fit energy well above 100 PeV [15]. This event could potentially be the first cosmogenic neutrino ever detected [16]. However, to make conclusions from this observation, one should take into account that events of similar energies have not been detected by other neutrino telescopes, despite larger exposure. These non-detections constrain strongly the isotropic neutrino flux, see e.g. [17, 18].
@@ II Data
For the present analysis, we use the Baikal-GVD sample of cascade events detected between Spring 2018 and Spring 2024. In this period, the detector was operated in several, increasingly larger configurations, with one or two new “clusters” of optical modules (OMs) added each spring. One cluster comprises 288 OMs arranged on eight vertical strings, with 36 OMs per string. The layout of the detector is presented in Fig. 1.
The inter-cluster spacing is varied in the range 250–300 m. Given the typical geometrical size of the cascade, the installation effectively works as a set of independent clusters in the cascade mode. Over the stated period, the number of clusters increased from three to twelve. The total exposure of the data set used here corresponds to years of one-cluster operation. The reader may find descriptions of the experiment, event reconstruction, analysis techniques and procedures e.g. in Refs. [19, 20, 21, 22, 23].
The search strategy for high-energy neutrino detection with Baikal-GVD was described in Ref. [20]. High-energy cascade events with the multiplicity of triggered optical modules (OMs) 19 and reconstructed energy TeV were selected, and additional cuts which suppress events from atmospheric muons were applied. The cut 19 was optimised for effective suppression of misreconstructed background events from atmospheric muon bundles. The number of these background events decrees rapidly with energy due to the steep energy spectrum of atmospheric muons.
In the present work, we select high-energy cascade events with 8 and impose the quality cuts described in Ref. [20], except for the cut introduced to suppress atmospheric muon bundles. Relaxing this latter cut and multiplicity of OMs allows us to increase the exposure considerably. Simulations indicate that the number of background events is about for all energies above 3 PeV, and above 10 PeV with these relaxed cuts.
@@ III Flux limits
where , is the effective area, the integration extends over the bins with in energy, over the celestial sphere in the solid angle and over the data collection time in . The effective area, measured in m2, is estimated by Monte-Carlo simulations and takes into account the efficiency of neutrino detection for a particular set of selection cuts. The resulting upper limits are listed in Table 1
@@ IV.2 Constraints on the cosmogenic neutrino flux models
Predictions of the selected models are shown in gray color in both Fig. 3 and Fig. 5. To obtain quantitative constraints on the scenarios, we make use of Eq. (1) and substitute there the theoretically predicted flux . In this way, we calculate, for each of the models, the expected number of events in the KM3-230213A 90% CL energy uncertainty band, in which the observed number of events is one. These theoretical numbers are presented in Table 2.
| model | |||
| Auger low 2023 [30] | 0.006 | 610 | |
| Auger high 2023 [30] | 0.34 | 11.5 | |
| Muzio maximal 2023 [31] | 7.0 | 0.56 | |
| Ehlert dip 2023 [13] | 4.1 | 0.96 | |
| KM3NeT 2025 extreme [16] | 0.98 | 3.9 |
Assuming that the model cosmogenic neutrino spectrum scales with the normalization factor , where for the theoretical model and the spectral shape does not change, we obtain 90% CL upper limits on from Eq. (1). We then make a temporary assumption that the single observed event was cosmogenic, and use the Poisson statistics [27] to get the best-fit value of and its 68% confidence interval. These numbers are also presented in Table 2. We see that the assumption of the cosmogenic origin of KM3-230213A is in tension with the lowest-flux Auger 2023 model, with the Poisson p-value corresponding to about discrepancy.
@@ Acknowledgements
Data availability
The effective area of the Baikal-GVD experiment is available in an ancillary file with this paper’s arXiv version [33].